OpenStudy (pinkkjazzy3):

Find the equivalent expression to the complex number (1+i)(1−i). a.0 B.2 c.1+i d.1-i

OpenStudy (sleepyjess):

(1+i)(1-i) 1*1=1 1*-i=-i i*1=i i*-i=\(-i^2\) \(-i^2=1\) 1-i+i+1 1+1 2